OpenStudy (anonymous):

what are the first five terms in the arithmetic sequence An=8n, stating with n=1

OpenStudy (solomonzelman):

\(\large\color{black}{ A_n=8\times n }\) \(\large\color{black}{ a_1=8\times 1=8 }\) \(\large\color{black}{ a_2=8\times 2=16 }\) \(\large\color{black}{ a_3=8\times 3=24 }\) So you can already tell it is going like, \(\large\color{black}{ 8,~16,~24~... }\)

OpenStudy (solomonzelman):

Can you see the pattern? Can you write the first 5 terms.